OpenStudy (anonymous):

suppose y varies directly with x, and y = 10 when x = -3. what direct variaton equation relates x and y? what is the value of y when x = -1 ?

OpenStudy (anonymous):

y/x = 10/(-3) y=10/(-3)x y=10/3 when x = -1 idk o.0

OpenStudy (anonymous):

me iether..